OpenStudy (anonymous):

which number equals the square root of 1377? 51 3 times the square root of 17 17 times the square root of 3 9 times the square root of 17

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

1377/3 = 459 459/3 = 153 153/3 = 51 51/3 = 17 So, the factors are 3^4 * 17. Square root of that must be 9 times square root of 17.

jhonyy9 (jhonyy9):

sqrt1377 =sqrt(81*17)=sqrt(9*9*17)=9sqrt17